OpenStudy (anonymous):

Need help! find the centroid of the region inside the surfaces z = x^2 + y^2 and x^2 + y^2 + z^2 = 2

OpenStudy (anonymous):

kinda confused any help would be great

OpenStudy (anonymous):

The centroid of the surface itself or the centroid of the solid bounded by the two equations?

OpenStudy (anonymous):

hmm im not really sure hold on here's the entire question: Let W be the region inside the surfaces z = x^2 + y^2 and x^2 + y^2 + z^2 = 2. Find the centroid of W

OpenStudy (anonymous):

so the centroid of the surface im guessing?

OpenStudy (anonymous):

Ok. First off, you need to find the Volume of the solid bounded by the equations, and to do that you have to find where they intersect.

OpenStudy (anonymous):

so would I plug in z=x^2+y^2 into the equation x^2+y^2+z^2=2

OpenStudy (anonymous):

You could, but that would be hard to solve. x^2+y^2+z^2=2 is the equation of a sphere centered at the origin with radius 2^(1/2). z = x^2 + y^2 Is the equation of an elliptical parabloid opening in the positive z direction. Their intersection would be at the plane z = 1.

OpenStudy (anonymous):

okay that makes sense

OpenStudy (anonymous):

What you want to do next is to find the limits of the triple integration that would give you the volume of the solid.

OpenStudy (anonymous):

okay so it would probally be easier to use spherical coordinates

OpenStudy (anonymous):

Cylindrical would be easier. The solid is bounded below by z= x^2 + y^2 = r^2 and above by z = 1. The projection of the solid onto the xy plane is the circle x^2 + y^2 = 1, which would give r a lower limit of 0 and an upper limit of 1, and theta a lower limit of 0 and an upper limit of 2pi.

OpenStudy (anonymous):

okay I understand the bounds but when you are trying to find the centroid what would you integrate?

OpenStudy (anonymous):

Ah wait I've made a mistake on the bounds of z. Its bounded above by the sphere z = (2-r^2)^(1/2)

OpenStudy (anonymous):

Well, to find the centroid of a region, you need the volume, which is just the triple integral of the limits. The centroid can then be found by the equations listed here: http://www.physicsforums.com/showthread.php?t=244944

OpenStudy (anonymous):

so you use the same bounds and integrate xdV, ydV, and zdV to get the 3 coordinates of the centroid

OpenStudy (anonymous):

Also note that once you find the Volume, you don't have to find the x and y coordinates of the centroid because by symmetry they are both zero. The solid is symmetrical across both the x and y axes. You can check this by doing the integrations as well. This just leaves the z coordinate of the centroid, so your final answer should be something along the lines of (0,0,z) where z is the z coordinate of the centroid.

OpenStudy (anonymous):

Yes, use the same bounds and integrate over zdV, but make sure to divide the final answer by the volume.

OpenStudy (anonymous):

\\[[V=\int\limits\limits_{0}^{2\pi}\int\limits\limits_{0}^{1}\int\limits\limits_{r^2}^{\sqrt{2-r^2}} dV\]

OpenStudy (anonymous):

im trying to set up the integral now but I think I am still confused on something So I set up the triple integral z dzdrdtheta using the bounds you stated above but then I have to divide by the volume of the sphere z^2+y^2+x^2=2 above the line z=1

OpenStudy (anonymous):

okay that makes more sense now that i see it

OpenStudy (anonymous):

Also note that dV is r dz dr dtheta, because when you convert to cylindrical you have to add the extra r in.

OpenStudy (anonymous):

oooo thank you for reminding me I always forget that r

OpenStudy (anonymous):

Alright I think i finally have this figured out thank you sooooo much for walking me through it sorry it took me so long but I think I understand how to do it now